### v1

# Quantum Optical Electron Pulse Shaper

Abstract
Coherent control of ultrafast quantum phenomena benefits from the pulse-shaping capabilities allowing to modulate the envelope and instantaneous phase of optical fields on femtosecond time scales. While such control is available for optical fields, an analogy of a pulse shaper for freely propagating electrons is lacking. In this study, we theoretically demonstrate a method that enables near arbitrary light-based shaping of electron wave packets in the time domain. The method is based on the quantum phase modulation of electron waves by coherent light with time-dependent frequency leading to generation of spectrally separated electron energy side bands with shaped time-energy profiles and envelopes. Our results show that few femtosecond time durations can be achieved without additional spectral broadening of the electron wave packet, allowing one to reach the combination of high time, spatial, and spectral resolutions in ultrafast imaging and diffraction experiments with pulsed electron beams.
preprint: APS/123-QED
I Introduction
Pulsed electron sources have been developed with the aim of extending the exceptional spatial resolution of electron microscopes to the time domain. Ultrafast electron microscopy and diffraction evolved in complex tools enabling observation of dynamical phenomena occurring on nanoscale, such as structural and phase transitions [1, 2], magnetic [3, 4] and charge carrier dynamics [5, 6, 7], atomic motion [8] and plasmonic [9] and phonon [10, 11] dynamics.
Pulsed electron beams for ultrafast electron microscopy are routinely generated via laser-stimulated photoemission. The energy bandwidth of the emitted electron wave packet is determined by the geometry and material properties of the emitter, the applied electric field, and the spectrum of the photoemission light. The nonzero energy spread combined with dispersive propagation of electrons lead to elongation of the generated pulses during their acceleration and propagation to the specimen [12]. As a result, the shortest pulse duration of electrons achieved in electron microscopes without additional compression is on the order of few hundreds of femtoseconds [13, 14, 15]. Such time resolution allows one to study picosecond processes, but the time scales associated, for example, with coherent lattice vibrations or coherent electron dynamics are not directly accessible without additional compression of the electrons.
Electron pulse compression schemes are based on time-correlated spectral broadening through interaction with time-dependent electromagnetic forces and subsequent compression due to dispersive propagation of the broadened electron pulse [16, 17, 18]. Such manipulation can occur in either the classical or the quantum-coherent regime and has been utilized for the compression of electron pulses to a few tens of femtoseconds durations when utilizing radio-frequency fields [19] or even attosecond durations when using optical fields [20, 21, 22, 23, 24, 25, 26]. However, experiments requiring both high spectral and temporal resolutions would greatly benefit from a technique that enables compression of electron pulses in the time domain while keeping the energy bandwith of the electrons reasonably narrow to maintain spatial and spectral resolution for imaging and diffraction experiments.
Although the phase and envelope of ultrashort optical pulses can be manipulated in the time domain by various shaping methods [27, 28], such complex control has not been available for pulsed electron beams. Optical pulse shaping techniques are typically based on separate control of the phase and amplitude of different spectral components of the pulse [28]. Their coherent superposition then determines the time profile of the pulse envelope. However, ultrashort electron pulses differ strongly from their optical counterparts due to the fact that their coherence time is much shorter than the Fourier-limited pulse duration [29]. In contrast to fully coherent optical waves generated by lasers, the electron wave packets from time-separated regions within the same pulse do not coherently add up to shape the envelope of the pulse.
Here we propose a quantum-optical electron pulse shaping concept, which is based on the interaction of a partially coherent electron pulse with shaped optical fields with time-varying frequency. Due to the energy-momentum dependence mismatch between photons and electrons in free space [30], an efficient energy modulation of the electrons requires breaking the spatial inversion symmetry of the optical fields. The light-electron interaction can be facilitated by a nearby plasmonic nanostructure [20, 31, 32, 21, 33, 34, 14], a metallic or dielectric membrane [25, 35, 36, 37, 24], an evanescent optical field [38, 39, 40, 41, 42, 43] or by the ponderomotive interaction with multiple optical beams in vacuum [22, 23, 44]. When the electrons interact with quasi-monochromatic light, they absorb and emit an integer number of photons. The resulting electron energy spectrum consists of sidebands separated by integer multiples of the laser photon energy $\hbar\omega$ [45]. Here we propose to utilize light pulse with time-dependent frequency $\omega(\tau)$. In such case, the distance between the neighboring photon side bands generated in the electron energy spectrum in any instant of time $\tau$ is given by the instantaneous photon energy $\hbar\omega(\tau)$ [46].
By tailoring the optical pulses used for coherent manipulation of the electron wave function, the longitudinal phase space of individual photon sidebands within the electron pulse can be precisely controlled and shaped, constrained only by the bandwidth of the optical pulses and the limits of optical pulse shaping. Combined with subsequent dispersive propagation of electrons to the specimen and with spectrally resolved detection, this technique can be utilized to generate close to arbitrary shape of the electron pulse in the longitudinal phase space. In this work we focus on application of this principle to compression of electron pulses and compensation of nonlinear chirp of the electrons.
II Electron Pulse Shaping Principle
Figure 1: (a) The duration of the electron pulse after the photoemission is determined by the time envelope of the the triggering optical pulse and the spectral width is dependent on the photoemission conditions. (b) After the emission, for distance $z_{\text{el}}$ the electron pulse is accelerated to a desired energy and is let to propagate dispersively to the site of the pulse-shaping interaction. During the process the electron pulse is elongated in time and acquired a positive chirp. At the interaction site the electron pulse traverses a dielectric membrane e.g., illuminated by an optical pulse with modulated instantaneous photon energy $\hbar\omega(\tau)$ and intensity envelope $I(\tau)$. (c) The absorptions and emission of photon quanta result in the generation of tailored energy sidebands, where the instantaneous energy separation between the bands is determined by $\hbar\omega(\tau)$ and the instantaneous quasi-probability $I(\tau)$. (d) After propagating for a sufficient distance, the quesi-probability side bands are reshaped and some of them experience temporal compression, the propagation distance denoted $z_{\text{prop}}$. The final time duration of the selected electron sub-pulse is denoted as $\tau_{\text{com}}$ and in dependence on the light field parameters, can be made shorter than the initial pulse duration by a factor of $M$.
The stimulated inelastic interaction between a periodic vector potential of light and the electron wave function can be understood in a semi-classical and nonrecoil approximations as a periodic phase modulation of the electron wave function [30]. The final wave function after the interaction can be described as a superposition of discrete momentum states separated by the momenta of the interacting photons. This principle can be generalized to the case of optical fields with time-dependent frequency, provided that the change in light frequency within the electron coherence time is negligible [46].
In the quantum optical electron pulse shaper proposed here, the electron pulse is generated by photoemission from a photocathode, accelerated by a static electric field, and propagates to the plane where pulse-shaping interaction is applied [see the layout in Fig. 1 (a)]. We assume that the duration of an electron pulse immediately after photoemission is determined by the time envelope of the optical pulse incident on the cathode in the electron source [12]. In contrast, the electron coherence time depends on the photoemission conditions. Without additional monochromatization, the typical electron energy spread in an ultrafast transmission electron microscope equipped with a field emission source is $\delta E=$0.3-0.7 eV [13, 47] with the associated coherence time of electrons $\tau_{\text{coh}}=\hbar/\delta E\approx$2-5 fs [12, 29].
In the interaction plane, the electron pulse interacts with a shaped optical pulse at a dielectric membrane [Fig. 1 (b)], ensuring efficient coupling of electrons and photons [48]. The electric field of the optical pulse generated in an optical pulse shaper [27, 28] can be described using an envelope and a time-dependent frequency $\omega(\tau)$. Due to the coherent phase modulation of the electron wave function at the instantaneous frequency of the shaped light pulse, the separation of the generated electron energy sidebands $\hbar\omega(\tau)$ becomes a function of time [Fig. 1 (c)]. Moreover, their instantaneous populations can be controlled by the envelope of the optical pulse. These two virtually independent quantities give us a tool for complex shaping of the electron wave in the longitudinal phase space. Combined with subsequent dispersive propagation between the pulse shaping plane and a sample, the electron pulse can be structured in time domain leading to a single compressed pulse or a periodic train of compressed pulses in one of the electron energy sidebands [Fig. 1 (d)], which can be isolated from the rest of the electron distribution by spectral filtering.
III Theoretical model
The semiclassical non-relativistic Hamiltonian of a charged particle interacting with an electromagnetic field with vector potential $\mathbf{A}$ is:
$\hat{H}=\frac{1}{2m_{e}}\left(\hat{\mathbf{p}}+e\mathbf{A}\right)^{2}=\hat{H}_{0}+\hat{H}_{\text{int}},$ (1)
where $e>0$ is the elementary charge and $\hat{\mathbf{p}}$ is the electron momentum operator. The free-particle Hamiltonian $\hat{H}_{0}=\hat{\textbf{p}}^{2}/2m_{e}$ describes the evolution of the wavepacket in free space, while the interaction Hamiltonian $\hat{H}_{\text{int}}$ describes the interaction with an electromagnetic field.
III.0.1 Before interaction
We describe the electron state in time-energy space, which is fully equivalent to the conventional coordinate-momentum representation. For a narrow spectral width, the relativistic energy-momentum relationship is approximately linear: $E\approx E_{0}+v_{0}(p-p_{0})$, where $E_{0}$ and $v_{0}$ are the relativistic energy and group velocity at the central momentum $p_{0}$, respectively. The temporal profile of the pulse is captured in the propagation-shifted time $\tau=t-z/v_{0}$, assuming that the electrons are propagating in the $z$ direction.
Before the interaction the pure state of the electron pulse is expressed by the wavefunction in the energy domain:
$\tilde{\psi}(E)=\mathcal{N}_{E}e^{\left[-\frac{(E-E_{0})^{2}}{2\sigma_{E}^{2}}\right]}e^{\left[-i\frac{\alpha(E-E_{0})^{2}}{2\hbar^{2}}\right]}e^{\left[-i\frac{\beta(E-E_{0})^{3}}{6\hbar^{3}}\right]},$ (2)
where $\mathcal{N}_{E}$ is a normalization constant, $\sigma_{E}=\delta E/\sqrt{4\ln(2)}$ the energy width, $\alpha$ is the group delay dispersion (GDD) coefficient and $\beta$ is the third order group dispersion coefficient (TOD). The time-domain representation is connected to the energy representation via Fourier transform $\psi(\tau)=\mathcal{F}\left\{\tilde{\psi}(E)\right\}$. To account for a partially coherent state, we introduce the density matrix $\rho(\tau,\tau^{*})$
$\rho(\tau,\tau^{*})=\sum_{\tau_{0}}w_{\tau_{0}}\psi(\tau-\tau_{0})\psi^{\dagger}(\tau^{*}-\tau_{0}),$ (3)
where $\tau^{*}$ is the second time coordinate of the matrix, $w_{\tau_{0}}$ is the probability of emission at time $\tau_{0}$ given by the expression
$w_{\tau_{0}}=\mathcal{N}_{w}\exp{\left[-\frac{8\ln{2}\tau_{0}^{2}}{\tau_{\text{pe}}^{2}}\right]},$ (4)
where $\mathcal{N}_{w}$ is a normalization constant, $\tau_{\text{pe}}$ is the FWHM duration of the photoemission optical pulse intensity envelope. Additionally, in the time representation of the wavefunction we only consider the envelope function of the plane wave, omitting the fast oscillation with frequency $E_{0}/\hbar$.
In phase space we represent the electron state via its Wigner function $W(E,\tau)$, derived from the density matrix $\rho(\tau,\tau^{*})$ via Wigner-Weyl transformation [49]. To model a dispersion-free state, we set $\alpha=\beta=0$ in Eq. (2), compute the partially incoherent density matrix $\rho_{0}(t,t^{*})$ via Eq. (3) and apply the Wigner transformation to obtain $W_{0}(E,t)$. This state has a probability density envelope FWHM duration of $\tau_{0}=\sqrt{\tau_{\text{coh}}^{2}+\tau_{\text{pe}}^{2}}$. For a dispersed state, we use nonzero $\alpha$ and $\beta$ to get $\rho_{1}(\tau,\tau^{*})$ and $W_{1}(E,\tau)$ with an extended FWHM duration $\tau_{\text{el}}=\sqrt{\tau_{0}^{2}+\tau^{2}_{\text{chirp}}}$, where $\tau_{\text{chirp}}$ accounts for dispersion induced elongation due to propagation between the photocathode and the interaction plane.
III.0.2 Interaction
The interaction with light occurs on much shorter time scales than the free space propagation, therefore it can be treated separately. In the non-recoil approximation, we calculate the effect of the interaction as a perturbation induced phase given by the integral of the Hamiltonian over the classical electron trajectories, described by the interaction time $\theta$ [50, 37, 21, 35]
$\phi(\textbf{r},t)=\left[-\frac{1}{\hbar}\int^{t}_{-\infty}\hat{H}_{\text{int}}(\textbf{r}+\textbf{v}_{0}(\theta-t),\theta)\,\mathrm{d}\theta\right],$ (5)
where $\mathbf{v}_{0}$ is the electron pulse group velocity vector. We define the evolution operator $\hat{U}(\textbf{r},t)=e^{i\phi(\textbf{r},t)}$. Applying the Coulomb gauge and the non-recoil approximation the interaction Hamiltonian becomes $H_{\text{int}}(\mathbf{r},t)=e\mathbf{p_{0}}\cdot\textbf{A}(\mathbf{r},t)/m_{e}$, where the ponderomotive term is neglected.
For a monochromatic electric field with frequency $\omega$, the vector potential is expressed in terms of the electric intensity $\mathbf{A}(\mathbf{r},t)=(-ic/\omega)\mathbf{E}(\mathbf{r},t)+c.c.$. A polychromatic optical incident field with spectrum ~$\tilde{\mathcal{E}}(\omega)=\sqrt{I(\omega)}e^{i\varphi(\omega)}$, where $I(\omega)$ is the spectral intensity and $\varphi(\omega)$ is the spectral phase, can be expressed as a superposition of the monochromatic components
$\textbf{E}_{i}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t+\mathbf{s_{i}}\cdot\mathbf{r}/c)},$ (6)
where $\mathbf{s}_{i}$ is the unit vector in the direction of propagation. In the time domain and space we can also express the incident electric field in terms of intensity envelope $I(\mathbf{r},t)$ and phase $\varphi(\mathbf{r},t)$
$\textbf{E}_{i}(\mathbf{r},t)\propto\sqrt{I(\mathbf{r},t)}\exp{\left[-i\varphi(\mathbf{r},t)\right]}.$ (7)
Expanding the phase into polynomial series of the wave argument $(t-\mathbf{s}\cdot\mathbf{r}/c)$ we obtain $\varphi(\textbf{r},t)=\omega_{0}(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)+a(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)^{2}+...$, where $\omega_{0}$ is the central angular frequency and $a$ is the linear chirp coefficient.
In the presence of a membrane or a nanostructure, the total electromagnetic field is
$\textbf{E}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}\mathbf{f}_{0}(\mathbf{r},\omega)e^{-i\omega t},$ (8)
where $\mathbf{f}_{0}(\mathbf{r},\omega)$ contains the spatial dependence of the field on frequency $\omega$ [48, 35].
For further evaluation, we consider 1D propagation of the electrons along $z$ with momentum vector $\mathbf{p}_{0}=(0,0,p_{0})$, group velocity vector $\mathbf{v}_{0}=(0,0,v_{0})$ and we only take into account the $z$ component of the electric field $f_{0z}$. Plugging into Eq. (8) into Eq. (5), we evaluate the evolution operator:
$\begin{split}U(z,t)=\\ \exp{\left\{-i2\text{Im}\int_{0}^{\infty}\mathrm{d}\omega\,g(\omega)\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t-z/v_{0})}\right\}},\end{split}$ (9)
where $g(\omega)$ is the interaction coupling strength for a given frequency
$g(x,y,\omega)=\frac{e}{\hbar\omega}\int_{-\infty}^{z}\mathrm{d}z^{\prime}\,f_{0z}(x,y,z^{\prime};\omega)e^{i\omega z^{\prime}/v_{0}}.$ (10)
Assuming that the coupling does not change significantly within the optical pulse bandwidth, we set $g(\omega)\approx g$. We transform to the shifted time coordinate $\tau=t-z/v_{0}$ and finally we obtain
$U(\tau)=\exp{\left\{-i2|g|\sqrt{I(\tau)}\sin{[\varphi(\tau)+\arg{g}]}\right\}}.$ (11)
In practice, we shift all the factors to $g$, so that $I(\tau)$ is normalized to 1 at the maximum. The density matrix describing the state of the electrons after the interaction is:
$\rho_{2}(\tau,\tau^{*})=U(\tau)\rho_{1}(\tau,\tau^{*})U^{\dagger}(\tau^{*}),$ (12)
and the corresponding Wigner function $W_{2}(E,t)$ is obtained via Wigner transformation.
III.0.3 After interaction
As a result of the interaction, sidebands with shaped time-dependent energy evolution are generated in the Wigner function $W_{2}(E,\tau)$, which are expected to be altered by dispersive propagation in vacuum for time $\theta_{\text{prop}}$. We emphasize that $\tau$ is the detection time within the electron pulse, whereas $\theta$ is the running evolution time. The evolution is described as coordinate transformation governed by the free-electron Hamiltonian $\hat{H}_{0}$ with the Hamilton equation of motion derived from its classical analogue $H_{0}$. The new coordinates after propagation time $\theta_{\text{prop}}$ are
$\tau^{\prime}-\tau^{\prime}_{0}=\tau-\tau_{0}+\frac{[v(E)-v_{0}]\theta_{\text{prop}}}{v_{0}},$ (13)
$E^{\prime}-E_{0}^{\prime}=E-E_{0},$ (14)
where $v[E(p)]=\partial E(p)/\partial p$ is the momentum-dependent group velocity. The transformed Wigner function is $W_{3}(E,\tau,\theta_{\text{prop}})=W_{2}(E^{\prime},\tau^{\prime})$.
IV Results
The pulse shaping method is examined in two model situations for lower kinetic energy electrons (5 keV) and higher kinetic energy electrons (30 keV). In the following section, we show the electron pulse represented by a non-negative spectrogram $S(E,\tau)$ defined as a convolution of the Wigner function with a Gaussian kernel
$S(E,\tau)=W(E,\tau)*\exp{[-\tau^{2}/\tau_{s}^{2}]},$ (15)
where $\tau_{s}\approx 5\,\mathrm{fs}$. This representation allows us to better visualize the energy band populations and smooths out the fast oscillating coherences between the bands [46].
IV.1 Chirp Inversion Induced Pulse Compression
We illustrate the temporal compression principle on the example of low energetic electrons with central kinetic energy of $5\,\mathrm{keV}$. Low-energy electrons, in general, undergo significant temporal elongation due to dispersion already after short distances, reaching durations several times longer than the initial length of the pulse. Also, the compression of the low energetic electrons can be conveniently achieved in relatively short propagation lengths.
We model the chirp-free electron pulse ($\alpha=0=\beta$ in Eq. (2)), with coherence time $\tau_{\text{coh}}=3.65\,\mathrm{fs}$, energy spread FWHM of $\delta E=0.5\,\mathrm{eV}$ and photoemission pulse duration $\tau_{\text{pe}}\approx 9.30\,\mathrm{fs}$, yielding an electron pulse with total FWHM duration of $\tau_{0}=9.9\,\mathrm{fs}$ [Fig. 2 (a)]. The central group velocity is $v_{0}=0.14c$.
We introduce second order group delay dispersion term only $\alpha=328\,\mathrm{fs^{2}}$, resulting in elongation of the pulse to $\tau_{\text{el}}=250\,\mathrm{fs}$ FWHM. The spectrogram of the chirped electron pulse is shown in [Fig. 2 (b)].
The spectral domain representation of the optical pulse we use for chirp inversion is explicitly
~$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}\right],$ (16)
where $\omega_{0}=1.8\,\mathrm{fs^{-1}}$ is the central photon frequency, the spectral width is matched to that of the electron wave packet $\sigma_{\omega}=\sigma_{E}/\hbar$ and the opposite GDD is used $\varphi^{\prime\prime}=-\alpha$. We consider constant coupling strength of $g=1.75$ to maximize the population of the chirp-inverted sideband.
The spectrogram [Fig. 2 (c)] of the electron pulse after the interaction consists of the ladder of energy bands spaced by integer multiples of the photon energies. This choice of instantaneous photon energy leads to the generation of a sideband with an inverted chirp with respect to the original electron wavepacket $E(\tau)+2\hbar\omega(\tau)$,centered around $E_{0}+2.4$ eV, corresponding to absorption of two photons. Around $20\%$ of the electron probability distribution is transferred to this band.
Under these conditions ($E_{0}\gg\hbar\omega$, $[v(E)-v_{0}]\propto E$), we can derive the required propagation time $\theta_{\text{com}}$, after which the side-band will compress. If the central group velocity of the second sideband is $v_{2\hbar\omega_{0}}=v(E_{0}+2\hbar\omega_{0})$, then the electron pulse needs to propagate for $z_{\text{com}}=\frac{m_{e}v_{2\hbar\omega}^{2}\alpha}{\hbar}\approx 21\,\mathrm{cm}$. We note this position as the compression point [Fig. 2 (d), (e)]. The side-band is compressed down to the initial duration of the pulse (numerically evaluated $\approx 9.9\,\mathrm{fs}$) reversing the elongation while maintaining the initial spectral width.
Figure 2: Spectrogram representation of the electron pulse (a) dispersion-free, with total pulse duration of 9.9 FWHM, (b) with group delay dispersion $\alpha$ and total duration of 250 fs FWHM. Spectrogram representation (c) of the electron side-bands after interaction with the optical field, (d) after propagation, (e) close-up of the compressed sideband spectrogram. We note that in the spectrogram representation the sidebands are elongated in the $\tau$-direction due to the convolution with Gaussian kernel.
IV.2 Nonlinear Electron Chirp Correction
Next we demonstrate the correction of higher order dispersion by introducing nonzero TOD to the electron wave. The parameters $E_{0}$, $\Delta E$ and $\alpha$ remain the same as in the previous case. The TOD is $\beta=256\,\mathrm{fs^{3}}$. The overall FWHM duration of the electron pulse is $\tau_{\text{el}}=242$ fs.
To compensate for $\beta$, we need to introduce TOD $\varphi^{\prime\prime\prime}$ into the optical pulse as well:
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\\ \exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right]\end{split}$ (17)
The curvature is compensated in the second side-band by a light pulse with the same GDD as previously and TOD $\varphi^{\prime\prime\prime}=-\beta/2=128\,\mathrm{fs^{3}}$.
Figure 3: Spectrogram of the electron pulse with 242 fs FWHM with linear and nonlinear chirp (a) after interaction with linearly chirped light (GDD only) (b) after interaction with light with GDD and TGD (c) GDD compensated after propagation for 21 cm, compression to 13 fs FWHM (d) after propagation for 21, GDD and TGD compensated, compressed down to 11 fs.
In Fig. 3 we show the spectrogram of an electron pulse with a combination of linear and nonlinear chirp. In [Fig. 3(a)] interaction with light field with $\varphi^{\prime\prime}=-\alpha$ and $\varphi^{\prime\prime\prime}=0$ is applied. In [Fig. 3(b)] interaction with light field with $\varphi^{\prime\prime}=-\alpha$ and $\varphi^{\prime\prime}=-\beta/2$ is applied. The difference between the two is subtle immediately after the interaction. After further propagation for 21 cm it is noticeable, that the sideband generated with linearly chirped light cannot compress ideally (c) due to the leftover curvature of the band, whereas introducing the $\varphi^{\prime\prime\prime}$ term leads to better compression (d).
IV.3 Short Pulse Gating
Until now, we have focused on modulation using long optical pulses that fully encompass the electron pulse. However, if the electron current is sufficiently high, it becomes possible to use optical pulses that are several times shorter than the electron pulse. This method requires shorter propagation distances for compression and is compatible with electron pulses of higher central energies. Additionally, it allows the final compressed sideband duration to be reduced below the photoemission duration, making the approach suitable even for pulses with long photoemission times.
We model the short pulse gating on an electron wavepacket with increased central kinetic energy of $30\,\mathrm{keV}$. The energy bandwidth $\delta E$ and the coherence time $\tau_{\text{coh}}$ remain unchanged. The initial duration of the electron pulse is increased to $\tau_{0}=50\,\mathrm{fs}$ [Fig 4(a)] and the duration after dispersive propagation remains $\tau_{\text{el}}=250\,\mathrm{fs}$ FWHM. The GDD coefficient is $\alpha=321\,\mathrm{fs^{2}}$, and the TDG coefficient is $\beta=0$ for simplicity.
While maintaining a constant spectral width of the optical pulse of $0.5\,\mathrm{eV}$, we set the GDD coefficient to $\varphi^{\prime\prime}=-\alpha/n$, where $n=11$. This is consistent with increasing the chirp $a$ while decreasing the pulse duration $\tau_{\text{opt}}$ approximately $n$-times. After the interaction a short, steeply chirped sideband centered around $1.2\,\mathrm{eV}$ is produced [Fig 4(b)]. To optimize the population of the first side-band we set $g=1$
The propagation distance required for the compression of the first sideband is then reduced to $z_{\text{com}}\approx\frac{m_{e}v^{2}\alpha}{\hbar(n-1)}$ while the time duration of the pulse after compression is reduced down to $\tau_{\text{com}}\approx\tau_{\text{0}}/(n-1)$. Specifically for $n=11$, where $4\%$ of the electrons from the initial pulse are transferred to the first sideband, the compression distance is $z_{\text{com}}\approx 32\,\mathrm{cm}$ and the duration of the compressed electron pulse is $\tau_{\text{com}}\approx 6\,\mathrm{fs}$ [Fig 4(c,d)].
Figure 4: (a) Spectrogram of the long initial pulse with FWHM duration 50 fs. Spectrograms (b) of the chirped electron pulse after interaction with a short optical pulse, (c) of the electron pulse state after propagation in free space for $\approx$ 30 cm, (d) close up of the compressed sideband with FWHM temporal duration of $\approx$ 10 fs. Spectrograms (e) of the electron pulse state after interaction with phase- and amplitude-modulated optical field, (f) after propagation for 21 cm, (g) close up of the compressed pulse train.
IV.4 Periodic Gating
The proposed scheme can however be used to produce tailored shape in the sideband, that will compress into a train of few-fs pulses in the compression point. While a single few-femtosecond laser pulse generates a single ultrashort electron sub-pulse, a longer optical pulse with periodic modulation in phase and amplitude can produce a train of ultrashort electron sub-pulses.
We start with 30 keV, 50 fs long electron pulse, stretched to 250 fs FWHM, as in the previous subsection. The optical pulse properties for this case can be introduced clearly in the time domain via the time-dependent frequency $\omega(t)$. As we state above, periodic frequency modulation is required
$\omega(t)=\omega_{0}\left[1+m\cos{\left(\Omega t\right)}\right]+2at,$ (18)
where $0<m\leq 1$ is the modulation depth and $\Omega$ is the modulation frequency, which should be a few times smaller than $\omega_{0}$ for this application. The chirp term is present to compensate for the zero-loss band chirp. If an optical field with cosine frequency modulation is used, the spacing between the sidebands is also periodically modulated. The optical field in the frequency domain can be written as
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\sum_{k=-1}^{k=1}\sum_{n=-\infty}^{\infty}(1/2)^{|k|}i^{n}J_{n}\left(\frac{m\omega_{0}}{\Omega}\right)\times\\ \exp\left[-\frac{(\omega-\omega_{0,kn})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0,kn})^{2}}{2}\right],\end{split}$ (19)
where $\omega_{0,kn}=\omega_{0}-(k+n)\Omega$ and $J_{n}$ is the Bessel function of the first kind.
Observing the first side-band, we note that every half period of the frequency oscillation the electrons have negative chirp and then in the next half period the electrons have positive chirp. The chirp can be precisely controlled by the modulation depth $m$ and frequency $\Omega$. In our case we choose $\Omega=\omega/20$ and $m=0.25$.
Ideally, only the parts of the sideband which have negative chirp must be populated. This is achieved by amplitude modulation of the optical field $\sqrt{I(t)}\propto\left[1+\cos{\left(\Omega t\right)}\right]/2$. Now only within the selected half-periods is the light field strong enough for a side-band to be generated [Fig. 4(e)]. This leads to the generation of negatively chirped sub-pulses, separated in time by $1/2\pi\Omega$, and with chirp coefficients $\mathrm{d}E/\mathrm{d}t=-\hbar m\omega_{0}\Omega$. The compression occurs at $z_{\text{com}}\approx\frac{m_{e}v^{2}\alpha}{\hbar(m\omega_{0}\Omega/2a-1)}\approx 21\,\mathrm{cm}$ with the duration of the compressed pulse of $\tau_{\text{com}}\approx 6\,\mathrm{fs}$ [Fig. 4(f,g)]. The structured side-band contains around $10\%$ of the initial electron distribution.
In the ideal case, a train of femtosecond pulses compressed several times below the initial duration of the pulse is generated. Considering realistic limitations, such as the finite bandwidth and the finite resolution of a standard light modulator, this application is still feasible, although minor distortions are to be expected. The temporal envelope of the side-band is expected to be shortened and the sideband modulation depth might be reduced (see Supplemental Material for details [51]).
V Discussion
The presented scheme is practically limited by three main factors. Firstly, the spectral and temporal properties of the optical pulses that can be realistically achieved with current shaping technologies. Secondly, we need to consider the phase-matching conditions of the chosen light-electron interaction over the broad spectrum of the optical pulse. Lastly, the available trajectory length within the electron microscope column determines the central energy of the electrons that will be used and the specific shape of the optical pulse.
State-of-the-art light shapers enable precise control over the spectral phase of laser pulses. Our method requires broadband optical pulses with an energy bandwidth of about 0.5 eV to allow straightforward shaping through the introduction of GDD and third-order dispersion TOD. More complex pulse shaping—such as that needed for generating periodic pulse trains—demands a light shaper with both high spectral resolution and sufficient bandwidth. For instance, a shaper with 300 pixels and a bandwidth of 0.7 eV centered around 1.2 eV is expected to produce a pulse closely matching the ideal case. For further details, see Supplementary Material.
In the main text, we approximate the coupling parameter $g(\omega)$ as constant over the bandwidth of the light pulse. In reality, the spectral dependence of $g(\omega)$ must be taken into account, and it may be necessary to adjust $I(\omega)$ to achieve the desired temporal profile of the generated sidebands. Furthermore, we do not address the effects of transverse momentum transfer during the interaction, which are expected to introduce additional nonlinear chirp due to the unequal trajectory lengths associated with different energies within the sidebands. However, we note that this effect can be mitigated by iteratively shaping the optical field to also compensate for electron delays arising from transverse effects.
Additional temporal broadening can arise from the fact that, in a focused beam, some electrons must traverse longer paths. This effect can be compensated by slightly defocusing the optical field at the membrane in the interaction plane. The phase curvature of the focused beam can then be used to counteract the trajectory length differences of the electrons.
VI Conclusion
We theoretically demonstrate a technique for versatile free-electron manipulation through quantum-coherent interaction with intensity- and phase-shaped optical fields. This approach effectively transfers the state of the art of light pulse shaping techniques to the pulsed electron case, enabling precise control over the electron time-energy characteristics—including spectral width, temporal duration, and both second and higher-order dispersion—in a broadly applicable manner. We demonstrate that the quantum optical electron pulse shaper may achieve electron pulse compression of electrons to few-femtosecond pulse durations without additional spectral broadening of electrons, making this approach ideally suited for ultrafast experiments requiring high time, spatial and spectral resolutions.
Acknowledgements.
The authors acknowledge funding from the Czech Science Foundation (project 22-13001K), Charles University (SVV-2023-260720, PRIMUS/19/SCI/05, GAUK 90424) and the European Union (ERC, eWaveShaper, 101039339). This work was supported by TERAFIT project No. CZ.02.01.01/00/22_008/0004594 funded by OP JAK, call Excellent Research.
Data Availability
The data supporting the findings of this study are openly available at [52].
Supplemental Material
Supplemental Material: Quantum Optical Electron Pulse Shaper
Neli Laštovičková Streshkova Martin Kozák
July 21, 2025
I Feasibility considerations
Since the required optical phase and amplitude modulation would be achieved by an optical modulator with finite resolution and from a finite bandwidth, we investigate the behavior of the pulse shaper under realistic conditions.
Under idealized conditions, the frequency spectrum of the optical field is continuous and infinitely broad. However, in reality, both the bandwidth and the spectral resolution of optical pulse shapers are finite and will limit the practical implementation and performance of the proposed electron pulse shaper. We consider a finite sampling of $N=100$ pixels, spanning a finite spectral window, typically more narrow than that of the optical pulse. In our example, we use a rectangular window with the width of $\Delta\omega=$ 0.7 eV, centered at 1.2 eV, which is experimentally achievable [53].
We discuss the cases - the optical pulse with GDD and TOD dispersion and the optical pulse with periodic amplitude and phase modulation combined with GDD. The respective ideal spectral representations are:
~$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right],$ (S1)
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\sum_{k=-1}^{k=1}\sum_{n=-\infty}^{\infty}(1/2)^{|k|}i^{n}J_{n}\left(\frac{m\omega_{0}}{\Omega}\right)\times\exp\left[-\frac{(\omega-\omega_{0,kn})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0,kn})^{2}}{2}\right]\end{split}$ (S2)
We show the comparison between the interaction of the electron pulse with a light field with an ideal, continuously modulated spectral intensity and a light field, which models a realistic optical pulse that is shaped by a light modulator.
To model the ”realistic” light pulse, we first limit the spectral width by a rectangular window, which is non-zero for the interval $(\omega_{0}-\Delta\omega/2;\omega_{0}+\Delta\omega/2)$. Then we divide this interval by the number of pixels $N$ and assign a discrete amplitude and phase value to each of the frequencies, effectively truncating and down-sampling the light pulse frequency spectrum and spectral phase. Formally, the integration is replaced by the discrete sum
$\textbf{E}_{i}(\mathbf{r},t)=\frac{\Delta\omega}{N}\sum_{n=1}^{N}\tilde{\mathcal{E}}(\omega_{n})e^{-i\omega_{n}(t+\mathbf{s_{i}}\cdot\mathbf{r}/c)},$ (S3)
where $\omega_{n}=\omega_{0}-\Delta\omega(1/2-n/N))$.
In the case of GDD and TOD correction, the comparison of the ideal and realistic intensity spectra is shown in Fig. S1. Large part of the spectrum around the central frequency remains practically unchanged. Because the central most intense part of the spectrum is also the part that mainly contributes to the generation of the second sideband, the result of the interaction with the realistically shaped optical field is almost identical to the one with ideally shaped optical field.
On the other hand, in the case of the periodic frequency and intensity modulation of the optical pulse, the spectrum is more complicated, consisting of many peaks (see Fig. S3, blue line). While the discretization to 100 pixels is not detrimental to the peak resolution, important features of the spectrum are not transmitted through the spectral window, resulting in changes in the spectrum of the optical pulse Fig. S3, orange line. After interaction with the realistic optical pulse intensity spectrum, the sub-pulses have narrower spectrum [see Fig. (S4)]. However, our data demonstrate that downsampling of the spectral intensity and phase of the optical pulse shaper are not detrimental for the operation of the proposed electron pulse shaper.
Figure 5: Comparison of the ideal and the down-sampled and frequency-restricted spectrum of light pulse with GDD and TOD.
Figure 6: Wigner function of the electron pulse after interaction with the (a) ideal optical field with GDD and TOD, (b) after compression of the second side-band, (c) after itneraction with realistically shaped optical field and after compression (d).
Figure 7: Comparison of the ideal and the down-sampled and frequency-restricted spectrum of light pulse with periodic phase and amplitude modulation.
Figure 8: Wigner function of the electron pulse interaction with the ideal optical field with periodic phase and frequency modulation, (a) before and (b) after compression of the first side-band, after interaction with realistically shaped optical field (c) before and (d) after compression.